OpenStudy (anonymous):

Hi do anyone know how to do Rational Expressions. m+6/m^2-3m+2m-1/m^2+5m-14

OpenStudy (anonymous):

is it :\[\frac{m+6}{m^2-3m} + \frac{2m-1}{m^2+5m-14}?\]

OpenStudy (anonymous):

m+6m2−3m+2 +2m−1m2+5m−14 there's two (2)s but that setup is right

OpenStudy (anonymous):

no fractions? lol I'm confused now

OpenStudy (anonymous):

lol yes fractions i just dont know how to put it on. its my first time on this. your set up is right. but you were missing a number 2 on m^2-3m+2 =)

OpenStudy (anonymous):

try to put brackets for those with fractions :)

OpenStudy (anonymous):

oh, so my equation was right? all I have to do is add a 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

=) your soo awesome for helping me double =D

OpenStudy (anonymous):

lol ^_^

OpenStudy (anonymous):

thank you )

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

alright, so after factoring the denominators we'll get:\[\frac{m+6}{(m-1)(m-2)} + \frac{2m-1}{(m+2)(m-7)}\] now get them to have the same denominators: \[\frac{(m+6)(m+2)(m-7) + (2m-1)(m-1)(m-2)}{(m+2)(m-2)(m-1)(m-7)}\]

OpenStudy (anonymous):

hmmm, there's something wrong

OpenStudy (anonymous):

are you sure of your question?

OpenStudy (anonymous):

LOL! WAIT!

OpenStudy (anonymous):

I got the mistake >_< sorry!

OpenStudy (anonymous):

it's supposed to be : \[\frac{(m+6)(m+7) + (2m-1)(m-1)}{(m-2)(m-1)(m+7)}\]

OpenStudy (anonymous):

so :

OpenStudy (anonymous):

\[= \frac{m^2+13m+42+2m^2-3m+1}{(m-1)(m-2)(m+7)}\]\[= \frac{3m^2+10m+43}{(m-1)(m-2)(m+7)}\]

OpenStudy (anonymous):

hmm, I'm not sure if we can simplify any further, correct me if I;m wrong please :)

OpenStudy (anonymous):

lol, yeah I got the question, but I'm not sure of my answer :)

OpenStudy (anonymous):

I am sure of my steps though :)

OpenStudy (anonymous):

=) nthank you no your fine im very confuse about this but i trust your answer

OpenStudy (anonymous):

thank youuu soooo much!!!

OpenStudy (anonymous):

all I did was factor and get them to have the same denominator, and you're welcome ^_^

OpenStudy (anonymous):

=) but why did you do those steps

OpenStudy (anonymous):

or how? so i can do them for more problems